# From Hypothesis to Publication: A Comprehensive Survey of AI-Driven Research Support Systems

### v1

@@ Appendix A Further Discussion
We posit that AI for research constitutes a subset of AI for science. While AI for research primarily focuses on supporting or automating the research process, it is not domain-specific and places greater emphasis on methodological advancements. In contrast, AI for science extends beyond the research process to include result-oriented discovery processes within specific domains, such as materials design, drug discovery, biology, and the solution of partial differential equations (Zheng et al., 2023; AI4Science and Quantum, 2023; Zhang et al., 2024d).
@@ B.1 Hypothesis Formulation
Existing paper recommendation systems predominantly rely on the metadata of existing papers to recommend related articles, often lacking specificity. By LLMs, dynamic user profiles can be constructed to provide personalized literature recommendations and enhance the richness of associated information for recommended articles, ultimately improving the user experience. In the Systematic Literature Review phase, outline generation frequently produces repetitive results with insufficient hierarchical structure. Furthermore, the full-text generation process is susceptible to hallucinations, a common issue observed in LLMs (Huang et al., 2023; Bolaños et al., 2024; Susnjak et al., 2024).
@@
Pre-trained models that rely on prompts encounter challenges in balancing novelty and feasibility during hypothesis generation. Further optimization is necessary to dynamically adjust the relative emphasis on novelty, feasibility, and validity in this process (Li et al., 2024c). Moreover, existing research on hypothesis generation frequently employs novelty and feasibility as evaluation metrics; however, these metrics are characterized by significant uncertainty.
@@ B.2 Hypothesis Validation
Existing approaches to scientific claim verification are largely restricted to specific domains, thereby limiting their practical applicability  (Vladika and Matthes, 2023). In the field of theorem proving, challenges arise due to data scarcity and the absence of standardized evaluation benchmarks  (Li et al., 2024e). Meanwhile, experiment verification faces significant limitations, as automatically generated experiments often lack methodological rigor, practical feasibility, and alignment with the original research objectives  (Lou et al., 2024).
@@ B.3 Manuscript Publication
Similar to systematic literature surveys, manuscript writing is also adversely affected by hallucination issues (Athaluri et al., 2023; Huang et al., 2023). Even when forced citation generation is employed, incorrect references may still be introduced (Aljamaan et al., 2024). Furthermore, the text generated by models requires meticulous examination by researchers to avoid ethical concerns, such as plagiarism risks (Salvagno et al., 2023). AI-generated manuscript reviews frequently provide vague suggestions and are susceptible to biases (Chamoun et al., 2024; Drori and Te’eni, 2024). Additionally, during meta-review generation, models are prone to being misled by erroneous information originating from the manuscript review process.
@@
Validation LAB-Bench Laurent et al. (2024) Experiment Validation 2400 Multiple-choice Question Answer Accuracy, Precision, Coverage
@@
Literature Review  (§2.1.2) [ AutoSurvey  Wang et al. (2024e) , OpenScholar  Asai et al. (2024) , HiReview  Hu et al. (2024b), PaperQA2  Skarlinski et al. (2024), Knowledge Navigator  Katz et al. (2024), CHIME  Hsu et al. (2024), STORM  Shao et al. (2024), LitLLM  Agarwal et al. (2024a), Bio-SIEVE  Robinson et al. (2023), KGSum  Wang et al. (2022a), Agarwal et al. (2024b), Lai et al. (2024), Susnjak et al. (2024), Bolaños et al. (2024), Block and Kuckertz (2024), Zhu et al. (2023) , Altmami and Menai (2022) , leaf, text width=50em ] ] ] [ Hypotheses Generation  (§2.2) [ Other Works [ MOOSE-Chem  Yang et al. (2024c), IdeaSynth  Pu et al. (2024), Si et al. (2024), Kumar et al. (2024), Girotra et al. (2023), Qi et al. (2023), Krenn et al. (2022), Henry and McInnes (2017) , leaf, text width=50em ] ] [ Input Data Quality [ SciPIP Wang et al. (2024d), COI Li et al. (2024a), SciAgents Ghafarollahi and Buehler (2024), DATAVOYAGER  Majumder et al. (2024a), SciMON Wang et al. (2024c), Buehler (2024), Liu et al. (2024a) , leaf, text width=50em ] ] [ Hypothesis Quality [ Dolphin Yuan et al. (2025), VIRSCI Su et al. (2024), Nova Hu et al. (2024a), AI Scientist Lu et al. (2024), SGA Ma et al. (2024), HypoGeniC  Zhou et al. (2024), ResearchAgent Baek et al. (2024), Acceleron Nigam et al. (2024), MOOSE  Yang et al. (2024b) , leaf, text width=50em ] ] ] ] [ Hypothesis Validation  (§3) [ Scientific Claim Verification  (§3.1) [ Claim [ HiSS Zhang and Gao (2023), SFAVEL Bazaga et al. (2024), ProToCo Zeng and Gao (2023), Wührl et al. (2024a) , leaf, text width=50em ] ] [ Evidence [ Glockner et al. (2024a), Vladika and Matthes (2024b), Vladika and Matthes (2024a), Wührl et al. (2024b), Pan et al. (2023b), Glockner et al. (2022), Wadden et al. (2022b) , leaf, text width=50em ] ] [ Verification [ ClaimVer Dammu et al. (2024), MAGIC Kao and Yen (2024), FactKG Kim et al. (2023), aedFaCT Altuncu et al. (2023), PROGRAMFC Pan et al. (2023a), Zhang et al. (2024b), Muharram and Purwarianti (2024), Eldifrawi et al. (2024), Cao et al. (2024b), Das et al. (2023), Wu et al. (2023) , leaf, text width=50em ] ] ] [ Theorem Proving  (§3.2) [ DeepSeek-Prover Xin et al. (2024), Lean-STaR Lin et al. (2024), Thor Jiang et al. (2022a), COPRA Thakur et al. (2024), Logo-power Wang et al. (2024b), Baldur First et al. (2023), Mustard Huang et al. (2024c), DT-Solver Wang et al. (2023b), HTPS Lample et al. (2022), GPT-f Polu and Sutskever (2020), Wang et al. (2024a), Song et al. (2024), Li et al. (2024e), Zhao et al. (2023b), Yang et al. (2023c), Jiang et al. (2023) , leaf, text width=62.5em ] ] [ Experiment Validation  (§3.3) [ MatPilot Ni et al. (2024), CRISPR-GPT Huang et al. (2024a), AgentHPO  Liu et al. (2024c), AutoML-GPT Zhang et al. (2023), MLCopilot Zhang et al. (2024a), Arlt et al. (2024), Kambhampati et al. (2024), Ruan et al. (2024b), Huang et al. (2024b), Boiko et al. (2023), Bran et al. (2024), Szymanski et al. (2023), Liu et al. (2023b), Manning et al. (2024), Mou et al. (2024) , leaf, text width=62.5em ] ] ] [ Manuscript
@@
Generation [ Cocoa Feng et al. (2024), Step-Back Tang et al. (2024a), data-to-paper Ifargan et al. (2024), ARIES D’Arcy et al. (2024b), OREO Li et al. (2022), R3 Du et al. (2022), Lai et al. (2024), Chamoun et al. (2024), Pividori and Greene (2024), Jourdan et al. (2023), Lin (2024) , leaf, text width=50em ] ] ] [ Peer Review  (§4.2) [ Paper Review
@@
Generation [ PeerArg Sukpanichnant et al. (2024) GLIMPSE Darrin et al. (2024), MetaWriter Sun et al. (2024), CGI2 Zeng et al. (2023), Li et al. (2024b), Santu et al. (2024), Kumar et al. (2023), Li et al. (2023) , leaf, text width=50em ] ] ] ] ]

### v2

@@ Appendix A Further Discussion
We posit that AI for research constitutes a subset of AI for science. While AI for research primarily focuses on supporting or automating the research process, it is not domain-specific and places greater emphasis on methodological advancements. In contrast, AI for science extends beyond the research process to include result-oriented discovery processes within specific domains, such as materials design, drug discovery, biology, and the solution of partial differential equations (Zheng et al., 2023b; AI4Science and Quantum, 2023; Zhang et al., 2024d).
@@
Discussion: Potential links between artificial intelligence systems and human research practice
  • In research paper recommendation, PaperWeaver (Lee et al., 2024) offers an interactive page that allows users to modify the topics they are interested in.
  • In systematic literature review,  Block and Kuckertz (2024) highlights the significant role of humans, including setting correct questions and individualized problem-solving and theorizing. Meanwhile,  Hsu et al. (2024) emphasizes manual correction during the outline generation process.
  • In hypothesis generation, AI engages more closely with human researchers, ranging from scenarios where humans provide the core ideas and AI contributes by iteratively refining them  (Pu et al., 2024), to more collaborative settings where humans and AI engage in dialogue to facilitate new scientific discoveries  (Ni et al., 2024; Liu et al., 2024b; Ye et al., 2024).
  • In scientific claim verification,  Altuncu et al. (2023); Das et al. (2023) highlight the critical role of experts in countering fake scientific news and advocate for the incorporation of expert opinions as a form of evidence.
  • In theorem proving,  Song et al. (2024) proposes leveraging LLMs as assistants to human researchers by generating suggested proof steps throughout the proving process.
  • In experiment Validation,  Ni et al. (2024) enhances the experimental setup through human-AI dialogue, whereas  Li et al. (2024d) incorporates human input and real-time adjustments during the execution phase to optimize experimental design.
  • In manuscript writing,  Ifargan et al. (2024); Feng et al. (2024); Du et al. (2022a) require human intervention to suggest improvements to AI-generated paragraphs and enhance their quality through interactive methods.
  • In peer review,  Kumar et al. (2023); Darrin et al. (2024) advocate for assigning the responsibility of generating meta-reviews to human researchers. The role of AI is to assist by identifying conflicts among reviewers’ opinions and supporting the chair in the scoring process, rather than independently assigning scores.
@@
Discussion: Some areas that have not been discussed
In addition to the eight areas discussed above, there are other lines of work that also aim to support scientific research, such as reading assistance (Kang et al., 2020; Head et al., 2021; Lo et al., 2023), which helps researchers read academic papers; literature processing [^1] https://sdproc.org, which handles documents in various formats to provide effective data for subsequent tasks; as well as code and data generation (Bauer et al., 2024; Zheng et al., 2023a), which serve as a foundation for experimental validation. However, as our focus is on the core process of scientific research, we have chosen to omit these aspects from the main text.
@@ B.1 Hypothesis Formulation
Existing paper recommendation tools predominantly rely on the metadata of existing publications to suggest related articles, which often results in a lack of user-specific targeting and insufficiently detailed presentation that hampers comprehension. Leveraging LLMs can facilitate the construction of dynamic user profiles, enabling personalized literature recommendations and enhancing the richness of the contextual information provided for each recommended article, ultimately improving the user experience. In the process of generating systematic literature reviews, our practical experience reveals that the outline generation tools often produces redundant results with insufficient hierarchical structure. Moreover, the full-text generation process is prone to hallucinations—for instance, statements may not correspond to the cited articles—a pervasive issue in large language models (Huang et al., 2023; Bolaños et al., 2024; Susnjak et al., 2024). This problem can be ameliorated by enhancing the foundational model capabilities or by incorporating citation tracing.
@@
Most existing tools generate hypotheses by designing prompts or constructing systematic frameworks, which heavily rely on the capabilities of pre-trained models. However, these methods struggle to balance the novelty, feasibility, and validity of the hypotheses (Li et al., 2024c).Furthermore, our investigation reveals that many current approaches adopt novelty and feasibility as evaluation metrics; these metrics are either difficult to quantify or require manual scoring, which can introduce bias. To date, there is no unified benchmark to compare the various methods, and we believe that future research should prioritize establishing a unified metric that objectively reflects the strengths and weaknesses of different approaches.
@@ B.2 Hypothesis Validation
Most existing scientific claim verification tools are largely confined to specific domains, exhibiting poor generalizability, which limits their practical applicability  (Vladika and Matthes, 2023). Theorem proving, the scarcity of relevant data adversely affects performance improvements through training , results across different proof assistants are not directly comparable, and the lack of standardized evaluation benchmarks presents numerous challenges. Moreover, current approaches remain predominantly in the research stage and lack practical tools that facilitate interaction with researchers  (Li et al., 2024e). Experiment Validation, as automatically generated experiments often suffer from a lack of methodological rigor, practical feasibility, and alignment with the original research objectives  (Lou et al., 2024). All these fields require rigorous logical reasoning, and I believe that the recent surge in advanced reasoning technologies could potentially address these issues.
@@ B.3 Manuscript Publication
Similar to systematic literature surveys, manuscript writing is also adversely affected by hallucination issues (Athaluri et al., 2023; Huang et al., 2023). Even when forced citation generation is employed, incorrect references may still be introduced (Aljamaan et al., 2024). Furthermore, the text generated by models requires meticulous examination by researchers to avoid ethical concerns, such as plagiarism risks (Salvagno et al., 2023). AI-generated manuscript reviews frequently provide vague suggestions and are susceptible to biases (Chamoun et al., 2024; Drori and Te’eni, 2024). Additionally, during meta-review generation, models can be misled by erroneous information arising from the manuscript review process (Kumar et al., 2023). To address these issues, it may be necessary for the industry to establish appropriate regulations or to employ AI-based methods for detecting AI-generated papers and reviews (Lin et al., 2023a).
Appendix C Ethical Considerations
AI has demonstrated significant potential in enhancing productivity by mitigating human limitations, thereby motivating increased investigation into its capacity to accelerate the research process  (Messeri and Crockett, 2024). Nevertheless, its integration into scientific workflows introduces a range of ethical concerns  (Fecher et al., 2023; Morris, 2023), including algorithmic biases, data privacy issues, risks of plagiarism, and the broader implications of AI-generated content for research communities. In this work, we examine these ethical challenges across the key stages of the research lifecycle: hypothesis formulation, validation, and publication.
During the hypothesis formulation stage, research paper recommendation systems and literature reviews are commonly employed; however, they often suffer from limitations that can lead to the formation of information bubbles and restrict exposure to diverse viewpoints. Furthermore, these systems tend to reinforce recognition disparities between prominent and lesser-known researchers and may inadvertently contribute to the dissemination of misinformation  (Polonioli, 2021; Bolaños et al., 2024). To address these biases, recommendation algorithms can be enhanced by emphasizing content-based rather than author-based recommendations and by incorporating robust evaluation mechanisms to strengthen the credibility of suggested materials.
In contrast, AI-driven hypothesis generation presents more pronounced ethical challenges. First, the attribution of intellectual property rights and authorship for AI-generated hypotheses remains ambiguous  (Majumder et al., 2024a). Additionally, the widespread generation of low-quality content poses a risk of diluting the integrity of the academic landscape (Hu et al., 2024a), while the potential misuse of such technologies for illicit purposes cannot be overlooked  (Si et al., 2024). Addressing these concerns necessitates the development of robust accountability frameworks, the assignment of clear responsibility for AI-generated outputs to researchers, and the establishment of appropriate legal and regulatory mechanisms.
During the hypothesis validation phase, automated systems for scientific fact-checking remain underdeveloped. This limitation may be exploited by malicious actors to create advanced misinformation generators capable of circumventing existing fact-checking tools  (Wadden et al., 2022b). Likewise, in the context of experimental validation, there is a risk of unethical or legally questionable experiments being designed  (Eger et al., 2025). These concerns underscore the need for continued research into model safety.
During the manuscript publication stage, several challenges remain. Text generated by AI models may carry a risk of plagiarism  (Salvagno et al., 2023; Gupta and Pruthi, 2025), while AI-assisted peer reviews often offer vague feedback and exhibit inherent biases  (Schintler et al., 2023; Drori and Te’eni, 2024; Pataranutaporn et al., 2025). To address these issues, the development of robust detection methods is essential. However, current detection tools are still in the early stages of maturity  (Gupta and Pruthi, 2025).
Appendix D Ability Comparison
An effective survey should not only summarize existing methods within a field but also provide comparative analyses of different approaches. However, the domain of AI for Research remains in its early stages, with many areas lacking standardized benchmarks and even established evaluation metrics. To facilitate a clearer understanding of the distinctions among various methods, we draw on existing literature (Kang et al., 2023; Bolaños et al., 2024; Luo et al., 2025; Vladika and Matthes, 2023; Yang et al., 2023c; Li and Ouyang, 2022, 2024; Lin et al., 2023a) and adopt attribute graphs to compare representative approaches within each subfield, as illustrated in table  §1 to table  §8.
Method Human-Computer Interaction LLM Required Information Return Information Relevance Source
ComLittee (Kang et al., 2023) - Authorship Graphs Meta data with relevant authors R, Co, Ci
ArZiGo (Pinedo et al., 2024) - User Interest Meta data R
PaperWeaver (Lee et al., 2024) Collected Papers Meta data with description R
Kang et al. (2022) - - Author’s social network relationships +Reference relationship Meta data with relevant authors R
Table 1: Research Paper Recommendation, we referred to  Kang et al. (2023) for comparing different methods, where R represents Paper recommender score, Co represents Co-author relationship, and Ci represents Cited author relationship.
Method Research Field Across Stages Human Interaction Task Input Output Evaluation Method
AutoSurvey (Wang et al., 2024e) Any - Outline Generation, +Full-text Generation Title & Full Content Literature Survey LLM & Human
CHIME (Hsu et al., 2024) Biomedicine - Outline Generation Title & Full Content Hierarchical Outline Automatic Metrics
Knowledge Navigator (Katz et al., 2024) Any - - Outline Generation Title & Full Content Hierarchical Outline LLM & Human
Relatedly (Palani et al., 2023) Any - - Full-text Generation Title & Related Work Literature Survey Human
STORM (Shao et al., 2024) Any - - Outline Generation, +Full-text Generation Title & Full Content Literature Survey LLM & Automatic Metrics
Table 2: Scientific Literature Review, we referred to Bolaños et al. (2024) and made modifications, thereby comparing different methods.
Method Research Field Across Stages Human Interaction Multi-agent Trained Model Online RAG Novelty Feasibility Validity
COI (Li et al., 2024a) Any - - -
Learn2Gen (Li et al., 2024c) Artification Intelligence - - -
MatPilot (Ni et al., 2024) Materials Science - - -
SciAgents (Ghafarollahi and Buehler, 2024) Any - - - -
SciMON (Wang et al., 2024c) Any - - - - - -
Table 3: Hypothesis Generation, we referred to Luo et al. (2025) and made modifications, thereby comparing different methods.
Method Input Document Retrieval Human Interaction Rationale Selection Evidence Format Output
MULTIVERS (Wadden et al., 2022b) Claim & scientific abstract Provided - Longformer Document Label & sentence-level rationales
SFAVEL (Bazaga et al., 2024) Claim Pre-trained Language Model - - knowledge graph Top-K Facts & Corresponding Relevance Scores
ProToCo (Zeng and Gao, 2023) Claim-Evidence Pair Provided - - Sentence Label
MAGIC (Kao and Yen, 2024) Claim Provided - Dense Passage Retriever Sentence Label
aedFaCT (Altuncu et al., 2023) News Article Google Search Human Document Evidence
Table 4: Scientific Claim Verification, we referred to Vladika and Matthes (2023) and made modifications, thereby comparing different methods.
Method Generation Based Stepwise Heuristic Search Informal or Formal Human-authored Realistic Proof
IBR (Qu et al., 2022) - Informal -
GPT-f (Polu and Sutskever, 2020) - Formal
DT-Solver (Wang et al., 2023b) Formal
POETRY (Wang et al., 2024a) - - Formal
Table 5: Theorem proving, we referred to Yang et al. (2023c) and made modifications, thereby comparing different methods.
Method Research Field Across Stages Human Interaction Multi-agent Task Input External tools
AutoML-GPT (Zhang et al., 2023) Artification Intelligence - - - Automated Machine Learning Task-oriented Prompts -
Chemcrow (Bran et al., 2024) Chemistry - - Chemical Task Task Description
DOLPHIN (Yuan et al., 2025) Any - Automated Scientific Research Idea
MechAgents (Ni and Buehler, 2023) Physics - - Mechanical Problem - -
Manning et al. (2024) Social Science - Simulating Human - -
Table 6: Experiment Validation: we use attribute diagrams to compare different schemes, and the table design refers to Hypothesis Generation.
Method Across Stages Human Interaction Task Input Evaluation Method
AI Scientist (Lu et al., 2024) - Full-text Generation Manuscript Template & Experimental Results & Hypothesis LLM
data-to-paper (Ifargan et al., 2024) Full-text Generation Experimental Results & Hypothesis -
ScholaCite (Martin-Boyle et al., 2024) - - Related Work Generation Title & Abstract & Citation Citation Graph Metrics
SciLit (Gu and Hahnloser, 2023) - Citation Generation Keywords Automatic Metrics
Gu and Hahnloser (2024) - - Citation Generation Citation Intent & Keywords Human
Table 7: Manuscript Writing, we referred to Li and Ouyang (2022, 2024) and made modifications, thereby comparing different methods.
Method Across Stages Human Interaction Paper Review Meta Review Multi-agent Trained Model Output
Gamma-Trans (Muangkammuen et al., 2022) - - - - Peer-review Score
MARG (D’Arcy et al., 2024a) - - - - Peer-review Comments
CycleResearcher (Weng et al., 2024) - - - Peer-review Comments & Score
PeerArg (Sukpanichnant et al., 2024) - - - - - Final Decision
GLIMPSE (Darrin et al., 2024) - - - - Summary of Peer-review
Table 8: Peer Review, we referred to Lin et al. (2023a) and made modifications, thereby comparing different methods.
@@ Appendix D Ability Comparison
LAB-Bench Laurent et al. (2024) Experiment Validation 2400 Multiple-choice Question Answer Accuracy, Precision, Coverage
PaperBench Starace et al. (2025) Experiment Validation 20 Paper, Additional Information Code Replication Score
@@
Validation ScienceAgentBench Chen et al. (2025) Experiment Validation 102 Task Instruction, Dataset Information +Expert-Provided Knowledge Program Valid Execution Rate, Success Rate, CodeBERTScore, API Cost
@@
SCHOLAWRITE Wang et al. (2025) Manuscript Writing 62,000 Before-text Writing Intention, After-text F1-score, Lexical Diversity, Topic Consistency, Intention Coverage
@@
AI-Researcher $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Literature Review  (§2.1.2) [ AutoSurvey  Wang et al. (2024e) , OpenScholar  Asai et al. (2024) , HiReview  Hu et al. (2024b), PaperQA2  Skarlinski et al. (2024), Knowledge Navigator  Katz et al. (2024), CHIME  Hsu et al. (2024), STORM  Shao et al. (2024), LitLLM  Agarwal et al. (2024a), Bio-SIEVE  Robinson et al. (2023), KGSum  Wang et al. (2022a), Agarwal et al. (2024b), Lai et al. (2024), Susnjak et al. (2024), Bolaños et al. (2024), Block and Kuckertz (2024), Zhu et al. (2023) , Altmami and Menai (2022) , leaf, text width=50em ] ] ] [ Hypotheses Generation  (§2.2) [ Other Works [ MOOSE-Chem  Yang et al. (2024c), IdeaSynth  Pu et al. (2024), Si et al. (2024), Kumar et al. (2024), Girotra et al. (2023), Qi et al. (2023), Krenn et al. (2022), Henry and McInnes (2017) , leaf, text width=50em ] ] [ Input Data Quality [ SciPIP Wang et al. (2024d), COI Li et al. (2024a), SciAgents Ghafarollahi and Buehler (2024), DATAVOYAGER  Majumder et al. (2024a), SciMON Wang et al. (2024c), Buehler (2024), Liu et al. (2024a) , leaf, text width=50em ] ] [ Hypothesis Quality [ Agentrxiv Schmidgall and Moor (2025) Dolphin Yuan et al. (2025), VIRSCI Su et al. (2024), Nova Hu et al. (2024a), AI Scientist Lu et al. (2024), SGA Ma et al. (2024), HypoGeniC  Zhou et al. (2024), ResearchAgent Baek et al. (2024), Acceleron Nigam et al. (2024), MOOSE  Yang et al. (2024b) , leaf, text width=50em ] ] ] ] [ Hypothesis Validation  (§3) [ Scientific Claim Verification  (§3.1) [ Claim [ HiSS Zhang and Gao (2023), SFAVEL Bazaga et al. (2024), ProToCo Zeng and Gao (2023), Wührl et al. (2024a) , leaf, text width=50em ] ] [ Evidence [ Glockner et al. (2024a), Vladika and Matthes (2024b), Vladika and Matthes (2024a), Wührl et al. (2024b), Pan et al. (2023b), Glockner et al. (2022), Wadden et al. (2022b) , leaf, text width=50em ] ] [ Verification [ ClaimVer Dammu et al. (2024), MAGIC Kao and Yen (2024), FactKG Kim et al. (2023), aedFaCT Altuncu et al. (2023), PROGRAMFC Pan et al. (2023a), Zhang et al. (2024b), Muharram and Purwarianti (2024), Eldifrawi et al. (2024), Cao et al. (2024b), Das et al. (2023), Wu et al. (2023) , leaf, text width=50em ] ] ] [ Theorem Proving  (§3.2) [ DeepSeek-Prover Xin et al. (2024), Lean-STaR Lin et al. (2024), Thor Jiang et al. (2022a), COPRA Thakur et al. (2024), Logo-power Wang et al. (2024b), Baldur First et al. (2023), Mustard Huang et al. (2024c), DT-Solver Wang et al. (2023b), HTPS Lample et al. (2022), GPT-f Polu and Sutskever (2020), Wang et al. (2024a), Song et al. (2024), Li et al. (2024e), Zhao et al. (2023b), Yang et al. (2023c), Jiang et al. (2023) , leaf, text width=62.5em ] ] [ Experiment Validation  (§3.3) [ MatPilot Ni et al. (2024), CRISPR-GPT Huang et al. (2024a), AgentHPO  Liu et al. (2024c), AutoML-GPT Zhang et al. (2023), MLCopilot Zhang et al. (2024a), Arlt et al. (2024), Kambhampati et al. (2024), Ruan et al. (2024b), Huang et al. (2024b), Boiko et al. (2023), Bran et al. (2024), Szymanski et al. (2023), Liu et al. (2023b), Manning et al. (2024), Mou et al. (2024) , leaf, text width=62.5em ] ] ] [ Manuscript
@@
Generation [ Cocoa Feng et al. (2024), Step-Back Tang et al. (2024a), data-to-paper Ifargan et al. (2024), ARIES D’Arcy et al. (2024b), OREO Li et al. (2022), R3 Du et al. (2022a), Lai et al. (2024), Chamoun et al. (2024), Pividori and Greene (2024), Jourdan et al. (2023), Lin (2024) , leaf, text width=50em ] ] ] [ Peer Review  (§4.2) [ Paper Review
@@
Generation [ Deeprview Zhu et al. (2025) PeerArg Sukpanichnant et al. (2024) GLIMPSE Darrin et al. (2024), MetaWriter Sun et al. (2024), CGI2 Zeng et al. (2023), Li et al. (2024b), Santu et al. (2024), Kumar et al. (2023), Li et al. (2023) , leaf, text width=50em ] ] ] ] ]